A new car leaves a parking lot with its odometer indicating 23 miles. Two hours later it returns with an odometer reading Which statement is true of the cars trip? of 123 mile Its average speed was 50 m.p.h, its displacement was 100 miles, and its average velocity was zero. Its average speed was zero, its displacement was zero, and its average velocity was 50 m.p.h. Its average speed was zero, its displacement was 100 mies, and its average velocty was 50 m.p.h. Its average speed was 50 m.p.h, its displacement was zero, and its average velocity was zero. O cating that a red sports car passed mile marker 3 on a highway with a speed limit of 60 m.p.h. Exactly five minut later the car passes the cop, who is sitting at mile marker number 9. Even though the speed limit, was the car speeding at any point between markers 3 and 9 cops radar indicates that the car is going the O Yes, the cars average speed was 72 m.p.h. O Yes, the car was going at least 90 m.p.h. No. There is no way to prove the driver broke the speed limit.
Answer #1

(1) Distance traveled by the car in 2.0 hour, d = 123 - 23 = 100.0 mile

So, average speed = distance traveled / time = 100.0 / 2.0 = 50.0 mph

However, the displacement of the car = 0 (since it returns to its original position)

So, average velocity = displacement/time = 0 / 2 = 0

Therefore, LAST OPTION is the correct answer.

(2) Distance between mile marker 3 and mile marker 9, d = 9 -3 = 6 mile

This distance is covered by the car in 5 minutes.

So, speed of the car during the journey from mile marker 3 to mile marker 9, v = (6/5) x 60 = 72 mile per hour

Therefore, the car is speeding up during the journey from mile marker 3 to mile marker 9 and its average speed was 72 mph.

Hence, first option is the correct answer.